The demand and supply equations is 2p2+q2=11 and p+2q=7. Find the equilibrium price and quantity, where p stands for price and q stands for quantity.


1
Expert's answer
2021-09-14T06:04:51-0400

"Demand\\
2p^2+q^2=11\\q^2=11-2p^2\\q=\sqrt{11-2p^2}\\supply\\p+2q=7\\2q=7-p\\q=3.5-\frac{1}{2}p"


"Qd=Qs\\\sqrt{11-2p^2}=3.5-\frac{1}{2}p\\
we \space get\space p=1\\q=3.5-0.5(1)\\=3"